James has $1.25 in nickels and dimes. He has 3 times as many nickels as dimes. Find the number of each type of coin.
I know the answer but I'm not sure how to make the expression so please help.

Answer by Alan3354(69443)

James has $1.25 in nickels and dimes. He has 3 times as many nickels as dimes. Find the number of each type of coin.
==============
The hard way:
5n + 10d = 125
n = 3d
etc
===================
Easier:
3n + 1d = 25
125/25 = 5 sets
--> 15 nickels & 5 dimes